# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 4.1 Learning Intersections of Halfspaces with Adversarial Label Noise
First, we will prove (i) and then we apply it to prove (ii). Note that the expectation $\operatorname*{\mathbf{E}}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]$ is minimized if $R$ puts all of its mass at the tails of the function $(\mathbf{v}\cdot x)^{2}$, i.e., $R(\mathbf{x})=\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})$, where $y_{\epsilon}$ is the smallest number such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})]\leq\epsilon$. Hence, we have that
@@
$=\frac{1}{\sqrt{2\pi}}y_{\epsilon}e^{-{y_{\epsilon}}/2}-\epsilon$ (4)
@@
where in (4) we simply used partial integration. In (5) we used the fact that the function $xe^{-x/2}$ is decreasing when $x$ is larger than $2$, note that from A.1 we have that $y_{\epsilon}\gtrsim\log(1/\epsilon)\geq 2$, since $\epsilon$ is assumed to be less than some sufficiently small constant.
@@
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}\cdot\mathbf{x}R(\mathbf{x})]^{2}\leq\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\lesssim\epsilon^{2}\log(1/\epsilon)\;.$
@@
We next show that there exists a direction within $W$ along which the second moment is small.
@@
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p(\epsilon/K)+\mathrm{OPT}\log(1/\mathrm{OPT})\;.$
@@
Denote by $\mu\coloneqq\operatorname*{\mathbf{E}}[\mathbf{x}\mid f(\mathbf{x})]$, by $\bar{\mathbf{x}}\coloneqq\mathbf{x}-\mathbf{\mu}$ and by $\mathbf{M}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\bar{\mathbf{x}}\bar{\mathbf{x}}^{\top}f(x)]$. For a defining vector $\mathbf{w}^{(i)}$, by B.4 we have that $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1](1-ce^{-(\max(0,t_{i}))^{2}/2}),$ for a sufficiently small universal constant $c>0$. Furthermore, we claim that there exists $i\in[K]$ for which the bias $t_{i}$ is at most $2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$. To see that, note that if for every $i\in[K]$ we have that $t_{i}>2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})},$ then by the union bound
@@
which is a contradiction. Hence, it holds that $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1](1-c\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K)$ for some $i\in[K]$.
Therefore, we have that
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}$ $=(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}-\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]+\frac{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]^{2}}{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}$
$\leq-c\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+\frac{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]^{2}}{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}\;.$
@@
Note that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]=\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]+\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}(S(\mathbf{x},y)-T(\mathbf{x},y))]$. By the assumption that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq\epsilon p(\epsilon/K)$ for all $i\in[K]$, and the fact that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\geq\operatorname*{\mathbf{Pr}}[y=1]-\mathrm{OPT}\geq\epsilon$, it holds that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\leq p(\epsilon/K)$.
@@
Moreover, note that $\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]\rvert=\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}R(\mathbf{x})]\rvert$, for $R(\mathbf{x})=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$. Therefore, since $R(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d}$ and $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$, by Claim 4.5 we have that $\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}(S(\mathbf{x},y)]\rvert\lesssim\mathrm{OPT}\sqrt{\log(1/\mathrm{OPT})}$. Consequently, since $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\geq\mathrm{OPT}$ we have that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$. Using similar arguments we have that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}T(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$.
Combining the above results gives us
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p(\epsilon/K)+\mathrm{OPT}\log(1/\mathrm{OPT})\;.$
@@
Let $\mathbf{v}\in\mathbb{R}^{d}$ be a unit vector, and denote by $R(\mathbf{x})=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$. Note that $R(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d}$, and moreover $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$. Hence, by Claim 4.5 we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}S(\mathbf{x},y)]\rvert=O(\mathrm{OPT}\log(1/\mathrm{OPT}))$. By using similar arguments we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}T(\mathbf{x},y)]\rvert=O(\log(1/\mathrm{OPT})\mathrm{OPT})$. Therefore, we have that
@@
Denote by $t$ the quantity $t\coloneqq C^{\prime}\log(1/\mathrm{OPT})\mathrm{OPT}$, for sufficiently large constant $C^{\prime}$. Combining the Claims 4.6 and 4.7 we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y]$ $\lesssim-\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K-t+p(\epsilon/K)$
@@
where we used the fact that one of $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1],\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]$ is greater than $1/2$ and the assumption that for all labels $i\in\{0,1\}$, $\operatorname*{\mathbf{Pr}}[y=i]\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$, for a sufficiently large constant $C$. Setting $p(\epsilon/K)\leq c\epsilon/K$ for a sufficiently small universal constant $c>0$, concludes the proof of Proposition 4.4. ∎

### v2

@@ 4.1 Learning Intersections of Halfspaces with Adversarial Label Noise
First, we will prove (i) and then we will apply it to prove (ii). Note that the expectation $\operatorname*{\mathbf{E}}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]$ is minimized if $R$ puts all of its mass at the tails of the function $(\mathbf{v}\cdot\mathbf{x})^{2}$, i.e., $R(\mathbf{x})=\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})$, where $y_{\epsilon}$ is the smallest number such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})]\leq\epsilon$. Hence, we have that
@@
$=\frac{1}{\sqrt{2\pi}}y_{\epsilon}e^{-{y_{\epsilon}}/2}+\epsilon$ (4)
@@
where in (4) we simply used integration by parts. In (5) we used the fact that the function $xe^{-x/2}$ is decreasing for $x\geq 2$. Note that, by A.1, we have $y_{\epsilon}\gtrsim\log(1/\epsilon)\geq 2$, and since $\epsilon$ is assumed to be sufficiently small.
@@
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}\cdot\mathbf{x}R(\mathbf{x})]^{2}\leq\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\lesssim\epsilon^{2}\log(1/\epsilon)\;,$
@@
We next show that there exists a direction within $W$ along which the second moment of $f(\mathbf{x})$ is small. More precisely, we have the following:
@@
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p\bigg{(}\frac{\epsilon}{K}\bigg{)}+\mathrm{OPT}\log(1/\mathrm{OPT})\;.$
@@
Denote by $\mu\coloneqq\operatorname*{\mathbf{E}}[\mathbf{x}f(\mathbf{x})]/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]$ and define $\bar{\mathbf{x}}\coloneqq\mathbf{x}-\mathbf{\mu}$. Let $\mathbf{M}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\bar{\mathbf{x}}\bar{\mathbf{x}}^{\top}f(x)]$. For any defining vector $\mathbf{w}^{(i)}$, by B.4 we have
$(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\big{(}1-ce^{-(\max(0,t_{i}))^{2}/2}\big{)}\;,$
for a sufficiently small universal constant $c>0$. Next, we claim that there exists some $i\in[K]$ for which the bias $t_{i}$ is at most $2\sqrt{\ln({K}/{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$. To see this, assume for contradiction that for every $i\in[K]$ we have $t_{i}>2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$ and note that if $t_{i}>2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$ then $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\leq 0]<\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K$. By applying the union bound, we obtain
@@
which is a contradiction. Hence, there must exist some index $i\in[K]$ with $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1](1-c\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K)$. Fix such an index $i\in[K]$. Then, by definition,
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\,\mathbf{w}^{(i)}=\;(\mathbf{w}^{(i)})^{\top}\mathbf{M}\,\mathbf{w}^{(i)}-\mathbf{Pr}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]+\frac{\left(\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}\mathbf{w}^{(i)}\cdot\mathbf{x}\,f(\mathbf{x})\Big{]}\right)^{2}}{\mathbf{Pr}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}.$
Substituting the bound on $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\,\mathbf{w}^{(i)}$, we obtain
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\,\mathbf{w}^{(i)}\leq-c\,\frac{\mathbf{Pr}[f(\mathbf{x})=1]\,\mathbf{Pr}[f(\mathbf{x})\neq 1]}{K}+\frac{\left(\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}\mathbf{w}^{(i)}\cdot\mathbf{x}\,f(\mathbf{x})\Big{]}\right)^{2}}{\mathbf{Pr}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}.$
@@
Note that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]=\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]+\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}(S(\mathbf{x},y)-T(\mathbf{x},y))]$. Recall that by assumption we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq\epsilon p(\epsilon/K)$ for all $i\in[K]$. Moreover, using the triangle inequality $\mathbf{Pr}[f(\mathbf{x})=1]\geq\mathbf{Pr}[y=1]-\mathbf{Pr}[f(\mathbf{x})\neq y]$ together with the assumption in Proposition 4.4 that $\mathbf{Pr}[y=1]\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$, it follows that $\mathbf{Pr}[f(\mathbf{x})=1]\geq\mathrm{OPT}+\epsilon$. Thus,
$\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\leq p(\epsilon/K)\;.$
@@
Next, observe that $\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]\rvert=\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}R(\mathbf{x})]\rvert$, where we define $R(\mathbf{x})=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$. Since, for all $\mathbf{x}\in\mathbb{R}^{d}$ it holds that $R(\mathbf{x})\in[0,1]$, and $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$, by Claim 4.5 we have that $\lvert\operatorname*{\mathbf{E}}[(\mathbf{w}^{(i)}\cdot\mathbf{x})S(\mathbf{x},y)]\rvert\lesssim\mathrm{OPT}\sqrt{\log(1/\mathrm{OPT})}$. Consequently, since $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\geq\mathrm{OPT}$, it follows that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$. A similar argument shows that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}T(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$. Combining these bounds, we obtain
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p\bigg{(}\frac{\epsilon}{K}\bigg{)}+\mathrm{OPT}\log(1/\mathrm{OPT})\;,$
@@
Let $\mathbf{v}\in\mathbb{R}^{d}$ be a unit vector, and denote by $R(\mathbf{x})=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$. Note that $R(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d}$, and moreover $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$. Hence, by Claim 4.5 we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}S(\mathbf{x},y)]\rvert=O(\mathrm{OPT}\log(1/\mathrm{OPT}))$. By using similar arguments we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}T(\mathbf{x},y)]\rvert=O(\log(1/\mathrm{OPT})\mathrm{OPT})$. Therefore, by triangle inequality, we have that
@@
Denote by $t$ the quantity $t\coloneqq C^{\prime}\mathrm{OPT}\log(1/\mathrm{OPT})$, for sufficiently large constant $C^{\prime}$. Combining Claims 4.6 and 4.7, we have that there exists $i\in[K]$ such that
$(\mathbf{w}^{(i)})^{\top}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y](\mathbf{w}^{(i)})$ $\lesssim-\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K-t+p(\epsilon/K)$
@@
where we used the fact that one of $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]$ or $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]$ must be at least $1/2$, and the assumption of Proposition 4.4 that for every label $i\in\{0,1\}$ it holds that $\operatorname*{\mathbf{Pr}}[y=i]\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$, for a sufficiently large constant $C$. Setting $p(\epsilon/K)\leq c\epsilon/K$ for a sufficiently small universal constant $c>0$, we ensure that the error term $p(\epsilon/K)$ is negligible compared to the main term. This concludes the proof of Proposition 4.4. ∎